3. HB Bookstore is currently located in Eyup. Nowadays, the owner is considering to open another branch in Beyazit. Suppose that the bookstore has $40 ordering cost and the lead-time is currently one week and will be same for the second location if it is opened. The owner states that there is a S0.30 weekly holding cost per item. Based on the market surveys the following demand forecasts are made for the following 6 weeks if the owner decides to open the second branch Week 1Week 2 Week 3Week 4 Week 5 10 20 Location Week 6 12 20 10 20 Be 30 30 20 It is also determined that if HB decides not to open the second branch, than 10% of the demand of the Eyup branch will be lost because of the inconvenience of the current location. The accountant of the bookstore says that HB makes $2 profit for each book it sells. If the second branch is opened, both branches will have independent inventory management. They are both going to utilize the current inventory policy, which is ordering units whenever R units are left and will target current 97% service level. Assume that the demand is normally distributed. What is the demand forecast of HBs Eyup location if the owner decides not to open the Beyazit branch? a. b. Determine the (R,Q) parameters for HBs Eyup branch if the second branch is not opened. c. Determine the average weekly inventory cost and average weekly ordering cost for HBs Eyup branch if the second branch is not opened. Determine the (R,Q) parameters for both of the branches if the second branch is opened. Determine the average weekly inventory cost and average weekly ordering cost for both of the branches if the second branch is opened. Do you think HB should open the new branch? Why? d. e. f.
Answer #1

HB Bookstore

Location

Week 1

Week 2

Week 3

Week 4

Week 5

Week 6

Eyup

10

12

14

12

10

8

Beyazit

20

20

30

30

20

20

a. if HB decides not to open the second branch, 10% of the demand of the Eyup branch will be lost

demand forecast for HB's EYUP location:

Location

Week 1

Week 2

Week 3

Week 4

Week 5

Week 6

Eyup

9

11

13

11

9

7

(rounding off)

b. EOQ, Q=sqrt(2*Co*D/Cc)

Annual Demand, D

515

(assuming 52 weeks in a year)

Ordering Cost, Co

$           40

Carrying Cost, Cc

$        0.30

EOQ, Q=

371

units

Reorder Point, R = d*L+z*?d*?L

Lead Time, L

1

week

Average Weekly demand, d

10

units

service level

0.97

z

1.88

std dev of demand, ?d

1.89

Reorder Point, R =

13

units

c. total ordering cost= (D/Q)*Co

$     55.58

average weekly ordering cost

$        1.07

(assuming 52 weeks in a year)

total carrying cost = (Q/2)*Cc

$     55.58

average weekly carrying cost

$        1.07

(assuming 52 weeks in a year)

d. if second branch is opened:

Location

Week 1

Week 2

Week 3

Week 4

Week 5

Week 6

Eyup

10

12

14

12

10

8

Beyazit

20

20

30

30

20

20

Eyup:

Average weekly demand, d

11

units

Annual Demand, D

572

(assuming 52 weeks in a year)

Ordering Cost, Co

$           40

Carrying Cost, Cc

$        0.30

EOQ, Q=

391

units

Lead time, L

1

week

Average Weekly demand, d

11

units

service level

0.97

z

1.88

std dev of demand, ?d

2.10

Reorder Point, R =

15

units

Beyazit:

Average weekly demand, d

23

units

Annual Demand, D

1213

(assuming 52 weeks in a year)

Ordering Cost, Co

$           40

Carrying Cost, Cc

$        0.30

EOQ, Q=

569

units

Lead time, L

1

week

Average Weekly demand, d

23

units

service level

0.97

z

1.88

std dev of demand, ?d

5.16

Reorder Point, R =

33

units

Formulae used:

1 HBBookstore 3 Location 4 Eyup 5 Beyazit Week 1 10 20 Week 2 12 20 Week 3 Week 4 Week5 Week 6 30 30 20 20 7 a. if HB decides35 d. if secondbranch is opened 36 Location 37 Eyup 38 Beyazit Week 1 Week 2 Week3 Week4 Week 5 Week 6 30 20 20 40 Eyup: 41 A